OpenStudy (ashy98):

The table below shows the squares of different numbers: Number (x) 2 -2 3 -3 Square of the number (y) 4 4 9 9 Part A: Does the table represent y as a function of x? Justify your answer. Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent?

OpenStudy (ashy98):

@ElonaSushchik

OpenStudy (anonymous):

A) y=f(x)=x^2 B)y=f(150)=20+3*150=20+450=470